# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v2

@@ 7 Discussion, Limitations, and Future Work
Our experiments have focused on binary classification models, which are known to perform well in supervised learning [16]. However, there is no constraint in applying T-Explainer to regression problems or more complex models, such as deep learning models. In particular, the T-Explainer naturally supports regression or multi-class classification with minor adaptations, which we are currently working on.
@@
Although T-Explainer has demonstrated competitive results when applied to tree-based models on synthetic data, tree-based architectures impose extra challenges on XAI. Tree classifiers have non-continuous architectures in which constant values are stored in leaf nodes, rendering gradient-based methods inappropriate. T-Explainer’s current version is not yet fully developed to support tree-based models. We are currently designing a mechanism that enables the computation of gradients in tree-based models, extending T-Explainer to operate in a more general context.
@@
The present implementation of the T-Explainer can handle categorical attributes through an approach that discretizes categorical data into intervals, using continuous perturbations, enabling the computation of partial derivatives. However, the approach requires a model retraining to fit the continuous intervals induced in the columns containing one-hot encoded categorical attributes. Retraining introduces additional computational complexity. Getting around the retraining issue is another improvement we are currently working on. One alternative approach we are considering in this context is to apply the target encoding transformer [11], which converts each nominal value of a categorical attribute into its corresponding expected value.
@@
Another aspect under improvement is the $\mathbf{h}$ optimization module. As far as we know, there is no closed solution to determine an optimum value to $\mathbf{h}$ in this context. During our investigations, we observed that the finite difference computation presents a certain instability for instances near the model’s decision boundaries due to discontinuities imposed by the decision boundaries. The literature on numerical methods brings a number of alternatives to deal with discontinuities when approximating derivatives through finite differences [62, 50]. We refer to methods that develop Generalized Finite Differences (GFD) among the alternative approaches [57, 13]. As part of the T-Explainer’s core, we are investigating alternatives to refine the $\mathbf{h}$ optimization process while maintaining computational efficiency.
@@
We conducted experiments and comparisons using tabular data to create classification models for simplicity and because it is a typical setup for machine learning problems in academia and industry [16]. We highlight that our experimental setup followed the evaluation practices in recent literature [10, 61, 12], focusing on Neural Networks as black boxes [37] and including datasets where leading XAI approaches have stability and applicability difficulties. However, we designed the T-Explainer considering more complex data representations, such as 2D images, videos, or semantic segmentation. We are extending our approach to encompass simplification masking approaches, a strategy widely used by many XAI methods [47, 39].
@@
A comprehensive assessment of an explanation method requires quantitative [15] and qualitative [65] measures. We prioritized quantitative metrics in this work, particularly those for evaluating consistency, continuity, and correctness [42]. We are committed to comprehensively validating the T-Explainer using metrics beyond those discussed here.
@@ 8 Conclusion
In this paper, we presented the T-Explainer, a Taylor expansion-based XAI technique. It is a deterministic local feature attribution method built upon a solid mathematical foundation that grants it relevant properties such as local accuracy, missingness, and consistency. Moreover, T-Explainer is designed to be a model-agnostic method able to generate explanations for a wide range of machine learning models, as it relies only on the models’ outputs without accounting for their internal structure or partial results. The experiments and comparisons demonstrate that T-Explainer is stable and accurate compared to model-specific techniques, thus making it a valuable explanation alternative.
@@ Acknowledgments
This work was supported in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior–Brasil (CAPES)–Finance Code 001, and in part by São Paulo Research Foundation (FAPESP) under Grant 2022/09091-8. The opinions, hypotheses, and conclusions or recommendations expressed in this material are the authors’ responsibility and do not necessarily reflect the views of the funding agencies.

### v3

@@ 7 Discussion, Limitations, and Future Work
Our experiments have focused on binary classification models, which are known to perform well in supervised learning [12]. However, there is no constraint in applying T-Explainer to regression problems or more complex models, such as deep learning models. The T-Explainer naturally supports regression or multiclass classification with minor adaptations, which we are currently working on.
@@
Although T-Explainer has demonstrated competitive results when applied to tree-based models on synthetic data, tree-based architectures impose extra challenges on XAI. Tree classifiers have non-continuous architectures in which constant values are stored in leaf nodes, rendering gradient-based methods inappropriate. The current version of T-Explainer is not yet fully developed to support tree-based models. We are currently designing a mechanism that enables the computation of gradients in tree-based models, extending T-Explainer to operate in a more general context.
@@
The present implementation of the T-Explainer can handle categorical attributes. It uses an approach that discretizes categorical data into intervals using continuous perturbations, enabling the computation of partial derivatives. However, the approach requires a model retraining to fit the continuous intervals induced in the columns containing one-hot encoded categorical attributes. Retraining introduces additional computational complexity. Getting around the retraining issue is another improvement that we are currently working on. An alternative approach we are considering in this context is to apply the target encoding transformer [9], which converts each nominal value of a categorical attribute into its corresponding expected value.
@@
Another aspect under improvement is the $\mathbf{h}$ optimization module. To our knowledge, there is no closed solution to determine an optimum value of $\mathbf{h}$ in this context. During our investigations, we observed that the finite difference computation presents a certain instability for instances near the model’s decision boundaries due to discontinuities imposed by the decision boundaries. The literature on numerical methods offers many alternatives to deal with discontinuities when approximating derivatives through finite differences [47, 38]. We refer to methods that develop Generalized Finite Differences (GFD) among the alternative approaches [43, 10]. As part of the T-Explainer’s core, we are investigating alternatives to refine the $\mathbf{h}$ optimization process while maintaining computational efficiency.
@@
We conducted experiments and comparisons using tabular data to create classification models for simplicity and because it is a typical setup for machine learning problems in academia and industry [12]. We highlight that our experimental setup followed the evaluation practices in the recent literature [8, 46], focusing on Neural Networks as black boxes and including datasets where leading XAI approaches have stability and applicability difficulties. However, we designed the T-Explainer considering more complex data representations, such as 2D images, videos, or semantic segmentation. We are extending our approach to include simplification masking approaches, a strategy widely used by many XAI methods [36, 30].
@@
The T-Explainer framework provides other metrics to evaluate feature importance in addition to those included in this publication, released as a Python module. We are actively extending this module to encompass more metrics and ensure their robustness, efficiency, and flexibility. A comprehensive evaluation of an explanation method requires quantitative [11] and qualitative [50] assessments. We prioritized quantitative metrics in this work, particularly those for evaluating consistency, continuity, and correctness [33]. We are committed to comprehensively validating the T-Explainer using metrics beyond those discussed here.
@@ 8 Conclusion
In this paper, we present the T-Explainer, a Taylor expansion-based XAI technique. It is a deterministic local feature attribution method built on a solid mathematical foundation that grants it relevant properties. Moreover, T-Explainer is designed to be a model-agnostic method capable of generating explanations for a wide range of machine learning models, as it relies only on the models’ outputs without accounting for their internal structure or partial results. The experiments and comparisons demonstrate that the T-Explainer is stable and accurate compared to model-specific techniques, thus making it a valuable explanation alternative.
@@ Acknowledgments
This work was partially supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior–Brasil (CAPES)–Finance Code 001, by CNPq Grant 307184/2021-8, and by the São Paulo Research Foundation (FAPESP) Grant 2022/09091-8. The views expressed in this material are the authors’ responsibility and do not necessarily reflect the views of the funding agencies.